OpenStudy (calculusxy):

MEDAL!!! A ball is thrown straight up. What will be the instantaneous velocity at the top of its path? What will be its acceleration at the top? Why are your answers different?

OpenStudy (calculusxy):

@jim_thompson5910

OpenStudy (anonymous):

Is it not that when it goes upward, at the top of its path, velocity =0?

jimthompson5910 (jim_thompson5910):

OOOPS is correct Gravity is pulling the ball down. The initial throw pushes the ball up and the velocity decreases as the ball climbs higher because it's fighting gravity. Once the ball gets to the peak, the velocity hits 0. The acceleration due to gravity is always the same no matter how fast the ball is moving.

OpenStudy (calculusxy):

What would be its acceleration?

jimthompson5910 (jim_thompson5910):

The acceleration due to gravity is always the same no matter how fast the ball is moving. acceleration of gravity = 9.8 m/s^2 = 32 ft/s^2 (both are approximate)

OpenStudy (calculusxy):

How did you calculate that 9.8m/s^2 = 38ft?s^2 ?

jimthompson5910 (jim_thompson5910):

the acceleration of gravity only changes when the height is drastically different technically there is change, but it's so small that it's hard to notice

jimthompson5910 (jim_thompson5910):

it's just something you look up in a table, book, or something. or you memorize it

OpenStudy (calculusxy):

I mean i know that 9.8m/s^2 is the instantaneous speed of a free-falling object. but how did you just calculate the acceleration at the top? acceleration = gravity x time ?

jimthompson5910 (jim_thompson5910):

the acceleration is equal to the acceleration of gravity

OpenStudy (calculusxy):

Honestly, i have no idea to what you just said.

jimthompson5910 (jim_thompson5910):

imagine the object in free fall every second, the speed increases by 9.8 m/s

OpenStudy (calculusxy):

okay

jimthompson5910 (jim_thompson5910):

let t be the time elapsed since dropping the object and letting it free fall (ignore air resistance) t = 0 speed = 0 m/s t = 1 speed = 9.8 m/s t = 2 speed = 19.6 m/s (9.8 times 2 = 19.6) t = 3 speed = 29.4 m/s (9.8 times 3 = 29.4) etc etc

OpenStudy (calculusxy):

yeah...

OpenStudy (calculusxy):

so what's going to happen next?

jimthompson5910 (jim_thompson5910):

No matter where the object is, it's feeling the same force on it. So it undergoes the same acceleration throughout its journey going up and then coming back down

OpenStudy (zzr0ck3r):

acceleration is a vector. i.e. it has direction.

OpenStudy (calculusxy):

but my question was how did you get this: "acceleration of gravity = 9.8 m/s^2 = 32 ft/s^2 (both are approximate)"

OpenStudy (zzr0ck3r):

Isac Newton

OpenStudy (zzr0ck3r):

I think...

jimthompson5910 (jim_thompson5910):

Galileo discovered it long before Newton, but yeah pretty much

OpenStudy (zzr0ck3r):

did he discover the speed or just the thing?

jimthompson5910 (jim_thompson5910):

those two numbers are constants that you memorize or look up

OpenStudy (calculusxy):

so the acceleration is 32?

jimthompson5910 (jim_thompson5910):

32 ft per s^2, yes

jimthompson5910 (jim_thompson5910):

just the speed/acceleration I think @zzr0ck3r I don't think he actually came up with unifying theories like Newton did

OpenStudy (zzr0ck3r):

And almost got killed as a result. Ahh humans...

OpenStudy (calculusxy):

so i found this website where it said that the acceleration can just be defined as "g", i believe:

OpenStudy (zzr0ck3r):

yes. gravity is constant and we normally just say g because we are estimating it when we give an actually number.

OpenStudy (zzr0ck3r):

like \(\pi\) and 3.14. They are not the same but we use 3.14 when we need a close number.

jimthompson5910 (jim_thompson5910):

g = 9.8 m/s^2 or g = 32 ft/s^2 both are the same. Saying 'g' is a shorthand way to say "acceleration of gravity on earth at sea level"

jimthompson5910 (jim_thompson5910):

When you hear things like "the rocket is going 3 g's (idk I made up the number)" it means "the rocket is accelerating 3 times the acceleration of gravity" so the rocket's acceleration would be 29 m/s^2

OpenStudy (calculusxy):

Thank you. I have another question.

OpenStudy (calculusxy):

If a salmon swims straight upward in the water fast enough to break through the surface at a speed of 5m/s, how high can it jump above the water?

jimthompson5910 (jim_thompson5910):

Acceleration of gravity: g = 9.8 m/s^2 Initial Velocity Vo = 5 m/s Plug those values into h = -(g/2)t^2 + Vo*t We don't know anything about the angle in which it jumps, so ignore that. Assume it jumps straight up. Find the vertex to get the answer.

OpenStudy (calculusxy):

h = (-9.8/2)5^2 + 5*2

jimthompson5910 (jim_thompson5910):

t is unknown for now

OpenStudy (calculusxy):

Yeah that's what I was guessing

jimthompson5910 (jim_thompson5910):

you should get h = -4.9t^2 + 5t find the vertex of this to figure out the highest point

OpenStudy (calculusxy):

How would I figured out the vertex?

OpenStudy (calculusxy):

*figure

jimthompson5910 (jim_thompson5910):

h = -4.9t^2 + 5t is in the form h = at^2 + bt + c a = -4.9 b = 5 c = 0 plug the values of 'a' and 'b' into -b/(2a) to figure out the time value at the peak once you get this value, plug it into h = -4.9t^2 + 5t to find h

OpenStudy (calculusxy):

so t would be -b/(2a) ?

jimthompson5910 (jim_thompson5910):

the t value at the peak, yes